Calculate the volume of ammonia gas (in mL) that will exactly neutralize 31.68 mL of a 0.215!M solution of sulfuric acid
1
Expert's answer
2020-12-21T03:48:02-0500

"2NH_3 + H_2SO_4 = (NH4)_2SO_4"

The above is the balanced equation for the reaction of ammonia gas with sulfuric acid


number of moles of sulfuric acid = molar concentration × volume = 0.215 × 31.68/1000 = 0.0068 mol


from the reaction above,

1 mole of sulfuric acid reacts with 2 moles of ammonia gas

thus, 0.0068 mol of sulfuric acid reacts with 2 (0.0068) mol = 0.0136 mol


1 mole of a gas = 22.4dm³

0.0136 mol = x dm³


x = "\dfrac{0.0136×22.4}{1}" = 0.305 dm³


0.305 dm³ = 0.305 × 1000 mL = 305 mL



Therefore, 305 mL of ammonia gas will exactly neutralize 31.68 mL.